A motorist traveled from city A to city B at an average speed of 50 mph. On the return trip, he left at 3:00pm and drove at an average rate of 60 mph, making the trip in 33 minutes less time. What is the distance from city A to city B?
Answer by stanbon(75887)

A motorist traveled from city A to city B at an average speed of 50 mph. On the return trip, he left at 3:00pm and drove at an average rate of 60 mph, making the trip in 33 minutes less time. What is the distance from city A to city B?
-----------------
A to B DATA:
Rate = 50 mph ; time = x hrs ; distance = rt = 50x miles
----------------------
B to A DATA:
Rate = 60 mph ; time = x-33/60 hrs ; distance = rt = 60(x-(33/60)) miles
------------------------
EQUATION:
distance = distance
50x = 60(x-(33/60))
50x = 60x - 33
10x = 33
x = 3.3 hrs
distance from A to B = 50(3.3) = 165 miles
====================
Cheers,
Stan H.